if a/b = c and x/a = c then a =
(A) x+b (B) ± xb (C)±√xb D) ±√x+b e) (x+b)^2

Answer by palanisamy(496)

given, a/b = c and x/a = c
=> a/b = x/a
a^2 = bx
a = ±√xb